### v1

# Differentiable Modal Synthesis for Physical Modeling of Planar String Sound and Motion Simulation

Abstract
While significant advancements have been made in music generation and differentiable sound synthesis within machine learning and computer audition, the simulation of instrument vibration guided by physical laws has been underexplored. To address this gap, we introduce a novel model for simulating the spatio-temporal motion of nonlinear strings, integrating modal synthesis and spectral modeling within a neural network framework. Our model leverages physical properties and fundamental frequencies as inputs, outputting string states across time and space that solve the partial differential equation characterizing the nonlinear string. Empirical evaluations demonstrate that the proposed architecture achieves superior accuracy in string motion simulation compared to existing baseline architectures. The code and demo are available online.  [^1] https://bit.ly/dmsp-notion-demo  [^2] https://huggingface.co/spaces/szin94/dmsp
1 Introduction
The investigation of wave propagation along strings, encompassing both theoretical and experimental dimensions, has persisted for well over a century [1, 2]. In the relentless pursuit of verisimilitude and expressive fidelity in simulating wave phenomena, numerous studies have been investigated to bridge the gap between theoretical underpinnings and empirical sound measurements [3]. Advancements leveraging computational power to mimic the intricate physical processes in musical instruments have given rise to numerical sound synthesis, now a cornerstone in the field of virtual sound synthesis [4, 5]. Schwarz [6] presents a systematic study of parametric and physical models for music audio synthesis. Parametric models include signal models, such as spectral modeling synthesis [7]. Physical models encompass techniques such as modal synthesis, digital waveguides [8, 9], or finite-difference time-domain (FDTD) methods [4, 10].
Over recent years, the advancement of hardware acceleration for artificial intelligence has enabled the emergence of numerous techniques for neural audio synthesis [11], including autoregressive generation [12], adversarial training [13] with phase coherence [14], and approximated physical models [15]. The concept of differentiable digital signal processing (DDSP) was first introduced by Engel et al. [16], aiming to incorporate a known signal model into neural networks to achieve a domain-appropriate inductive bias. While the DDSP model can be considered a differentiable version of spectral modeling synthesis, a wide variety of works have explored the differentiable implementation of other audio signal processing methods, such as the subtractive method [17], waveshaping [18], and frequency modulation [19, 20]. Subsequent research has demonstrated various applications of DDSP, including music performance synthesis [21], speech synthesis and voice conversion [22, 23], and sound effect generation [24]. Renault et al. [25] extended DDSP to create a polyphonic synthesizer, explicitly modeling properties specific to piano strings, such as inharmonicity and detuning induced by string stiffness, based on a parametric model of these phenomena [26]. This model efficiently synthesizes piano sound from MIDI input, achieving a high mean opinion score on naturalness. However, it still shows room for improvement compared to sampling-based methods [27] and physical modeling methods [28].
Despite the growing recognition of DDSP as a promising sound synthesis methodology, its extension to physical modeling remains underexplored. In the context of musical sound synthesis, some attempts have been made to synthesize rigid-body contact sounds of arbitrary shapes using neural networks supervised by finite-element method (FEM) solvers. Jin et al. [29, 30] propose a neural network that predicts contact sounds from voxelized objects, inspired by the modal technique that synthesizes sound using eigenvalues and eigenvectors. Diaz et al. [31] leverage a differentiable infinite impulse response filter to synthesize contact sounds from rasterized occupancy grids in an end-to-end manner. These methods can interactively synthesize sounds for various contact conditions and materials with notable efficiency, circumventing the need for an offline optimization process typical of modal techniques. However, these methods, which resort to the FEM solver, are vulnerable in modeling the dynamic behavior (e.g., glissando, vibrato) or in simulating the motion of the object. This vulnerability demands the need for further research to enhance the capability of neural networks in physical modeling for more dynamic and complex sound synthesis scenarios.
In this regard, we propose a novel model for simulating the spatio-temporal motion of nonlinear strings, integrating modal synthesis and spectral modeling within a neural network framework. The proposed model leverages physical properties and fundamental frequencies as inputs, outputting string states across time and space that solve the partial differential equation (PDE) characterizing the nonlinear string. Empirical evaluations demonstrate that the proposed architecture achieves superior accuracy in string motion simulation compared to the baseline architectures. The main contributions are as follows:
  • We present differentiable modal synthesis for physical modeling (DMSP) that simulates dynamic nonlinear string motion by synthesizing sound using the physical properties of the string.
  • To the best of our knowledge, this is the first approach that is differentiable and can synthesize the motion and the sound of a musical string with a dynamic control over the pitch and the material properties.
  • We provide an extensive empirical evaluation demonstrating the importance of modal decomposition and the proper choice of loss function.
Figure 1: System overview. The DMSP model encodes the physical properties of a string (e.g., tension, stiffness, damping, and initial conditions) to estimate the displacement of the string plucked at pitch $f_{0}$ at a given time $t\in[0,\infty)$ and position $x\in\Omega$. By concatenating the DMSP outputs over the domain $(x,t)\in\Omega\times[0,\infty)$, the simulated motion of the string can be visualized. Sampling the outputs at a particular position $x$ allows one to hear the synthesized string sound, akin to listening with a stethoscope at the pickup position.
2 Background
2.1 Physical Modeling of Musical String Instrument
Linear Damped Stiff String. The string model discussed in this paper is a damped nonlinear stiff string. To introduce the nonlinear string, we first formulate the governing equations for the damped linear stiff string system and derive the corresponding modal solution.
$\partial_{tt}u=\gamma^{2}\partial_{xx}u-\kappa^{2}\partial_{xxxx}u-2\sigma_{0}\partial_{t}u$ (1)
The linear string of its length $L$, vibrating with wave speed $\gamma$, stiffness $\kappa$, and frequency-independent damping factor $\sigma_{0}$, is described by Equation 1. Given the initial conditions (IC) for $x\in\Omega=[-L/2,+L/2]$ as $u(x,0)=u_{0}(x)$ and $\partial_{t}u(x,0)=0$, and appropriate boundary conditions (BC), the corresponding solution $u(x,t)$ represents the motion of a damped linear stiff string. Particularly, for a clamped boundary condition, i.e., $u(\pm L/2,t)=\partial_{x}u(\pm L/2,t)=0$ for all $t\in[0,\infty)$, an analytic solution can be obtained as follows.
$u(x,t)$ $=\sum_{n=1}^{\infty}X_{n}(x)T_{n}(t)$ (2a)
$X_{n}(x)$ $=c_{1}\left(\sin\mu_{n}x-\frac{\sin\mu_{n}L/2}{\sinh\nu_{n}L/2}\sinh\nu_{n}x\right)+c_{2}\left(\cos\mu_{n}x-\frac{\cos\mu_{n}L/2}{\cosh\nu_{n}L/2}\cosh\nu_{n}x\right)$ (2b)
$T_{n}(t)$ $=e^{-\sigma_{0}t}\cos\underbrace{\sqrt{\mu_{n}^{4}\kappa^{2}+\mu_{n}^{2}\gamma^{2}-\sigma_{0}^{2}}}_{\omega_{n}}t$ (2c)
The derivation of Equation 2 can be found in Appendix C. The allowed values of $\mu_{n}$ and $\nu_{n}$ are determined by the boundary conditions, while the coefficients $c_{1}$ and $c_{2}$ are determined using the initial condition $\sum_{n=1}^{\infty}X_{n}(x)=u_{0}(x)$. Determining these values typically requires an offline numerical solving process, where the obtained values can be stored in memory for real-time computation of $u(x,t)$. The stiffness modeled by the 4th order derivative induces a hyperbolic solution, resulting in non-integer multiples of the mode frequencies $\omega_{n}$, which leads to physical inharmonicity. The damping factor $\sigma_{0}$ causes an exponential decay in the temporal amplitude.
Figure 2: The planar string system.
Nonlinear Damped Stiff String. The generalization of the linear wave Equation 1 to nonlinear string vibrations was first introduced by Kirchhoff [32] and Carrier [33]. The Kirchhoff–Carrier system models elastic strings in two dimensions, and when extended so that transverse and longitudinal motions are coupled, phantom partials can be exhibited, resulting in a richer timbre [34]. A model of such planar string vibration is as follows [35, 36].
$\partial_{tt}u$ $=\gamma^{2}\partial_{xx}u-\gamma^{2}\frac{\alpha^{2}-1}{2}\partial_{x}\left(q^{3}+2pq\right)-\kappa^{2}\partial_{xxxx}u-2\sigma_{0}\partial_{t}u+2\sigma_{1}\partial_{t}\partial_{xx}u$ (3a)
$\partial_{tt}\zeta$ $=\gamma^{2}\alpha^{2}\partial_{xx}\zeta-\gamma^{2}\frac{\alpha^{2}-1}{2}\partial_{x}\left(q^{2}\right)-2\sigma_{0}\partial_{t}\zeta+2\sigma_{1}\partial_{t}\partial_{xx}\zeta$ (3b)
Here, $u(x,t)$ and $\zeta(x,t)$ represent the transverse and longitudinal displacements of a string, respectively, for all $(x,t)\in\Omega\times[0,\infty)$. $q=\partial_{x}u$ and $p=\partial_{x}\zeta$ serves as the auxiliary coupling system, as $\partial_{t}q=\partial_{x}\partial_{t}u$ and $\partial_{t}p=\partial_{x}\partial_{t}\zeta$ [37]. A more detailed background on the derivation of Equation 3 can be found in Appendix B. The boundary condition, which may vary depending on the string being modeled, is chosen to be that of the clamped boundary condition:
$u(x,t)=\partial_{x}u(x,t)=0,\quad\forall(x,t)\in\partial\Omega\times[0,\infty).$ (4)
Given an initial condition $u_{0}:=u(x,0)$ defined on $x\in\Omega$, the solution $u(x,t)$ associated with the condition of Equation 4 simulates the motion of the string for physical modeling and sound synthesis of string instruments. Due to the coupling between Equation 3a and Equation 3b, the obtained solution exhibits features found in elastic strings, such as pitch glide and phantom partials. These features become more pronounced for larger displacements and are difficult to approach separately as in the linear case. The solution can be approximated through various physical modeling techniques such as finite-difference time-domain [38], digital waveguides [39], or functional transformation method [40].
Finite-difference Time-domain. One straightforward approach to tackling nonlinear PDEs such as Equation 3 would be employing finite difference approximation. This method, commonly known as finite-difference time-domain (FDTD), has a long and distinguished history and is widely accepted in fields such as fluid dynamics [41] and electromagnetics [42]. FDTD is particularly effective in solving nonlinear, multidimensional, and dynamic systems. The extensive literature on its applications encompasses a diverse range of domains, including musical acoustics [10]. A recent contribution by Lee et al. [43] introduces StringFDTD-Torch, an FDTD simulator tailored for modeling planar damped stiff strings akin to Equation 3. Leveraging PyTorch C++ extension, StringFDTD-Torch facilitates FDTD computations on both CPUs and GPUs. However, its current iteration lacks support for gradient backpropagation through the FDTD module, leaving room for enhancement, particularly in optimizing the gradient computation process, which is hindered by the substantial number of temporal recursions involved (evident by the large $N_{t}$ in $\mathcal{O}(N_{x}N_{t})$ of Table 1).
Modal Synthesis. As a more efficient approach to solve the nonlinear wave equations, the modal synthesis [44, 45, 46] decomposes the complex dynamics into contributions from a set of modes, whose spatial bases are eigenfunctions of the pertinent problem. Each mode exhibits distinct oscillations at complex frequencies, contingent upon the boundary conditions. For problems with real-valued parameters, these complex frequencies occur in conjugate pairs, and the "mode" is thus defined as the pair of such eigenfunctions and frequencies [10]. Modal synthesis involves two primary steps. Initially, in an offline phase (also labeled as ‘Pre-computation’ in Table 1), modal shapes and frequencies are discerned from the PDE system, considering both boundary and initial conditions. This information is encapsulated in what is known as a shape matrix. Subsequently, the solution is derived by combining the modal functions, each progressing at its natural frequency. Pre-computation requires $\mathcal{O}(N_{m})$ of recursions because $N_{m}$ shape matrices need to be computed, but a typical $N_{m}$ is typically hundreds to thousands of times less for $N_{t}$. Modal synthesis after this off-line process is very efficient as it allows us to obtain a solution for a given $x$ and $t$ without any recursion, but it is clear that the range of solutions that can be covered is bounded in that it relies on the ansatz $u(x,t)=X(x)T(t)$ for the separation of variables.
2.2 Differentiable Digital Signal Processing
In the field of neural networks, numerous audio researchers have been engaged in the development of techniques that leverage the ease of automatic gradient backpropagation in neural networks for the purpose of audio parameter estimation. To address the challenge of estimating the latent parameters of a sound, some approaches implement the synthesis part as-is using automatic differentiation package [43, 20] so that the gradient can back-propagate through it to update the parameters directly, while the majority of approaches train neural networks to estimate the parameters in an auto-encoder framework [16, 47, 22]. As one of the most seminal studies of the latter approach, DDSP is widely used to efficiently synthesize nonlinear and dynamic sounds. Based on the spectral modeling synthesis framework, the time-domain signal is modeled via short-time Fourier transforms (STFTs) divided into deterministic (harmonic) and stochastic (noisy) parts to synthesize the sound. As the causality of STFT frames is modeled through gated recurrent units (GRUs), DDSP requires as many recursions as $N_{r}$, the number of frames. The harmonics of a DDSP are given by an integer multiple of the fundamental frequency ($f_{0}$), and the noise is synthesized from filtered noise. These DDSPs, while still a remarkable advancement, have strong structural constraints on the deterministic part to capture enough perceptually rich tones such as inharmonicity due to stiffness or phantom partials due to nonlinearity. A study by Renault et al. [25] also points this out, and uses the parametric model [26] for inharmonicity and detune, but there is room for improvement as it is an approximation model that relies on instrument-specific modifiers rather than reflecting stiffness physics.
Table 1: Comparison between methods.
Method System taxonomies Computational complexity
Physical Nonlinear Differentiable Pre-computation Inference
Modal $\mathcal{O}(N_{m})$ $\mathcal{O}(1)$
FDTD - $\mathcal{O}(N_{x}N_{t})$
DDSP - $\mathcal{O}(N_{r})$
DMSP-Hybrid $\mathcal{O}(N_{m})$ $\mathcal{O}(1)$
DMSP - $\mathcal{O}(1)$
3 Differentiable Modal Synthesis for Physical Modeling (DMSP)
This section introduces a novel differentiable nonlinear string sound synthesizer. Table 1 provides a summary of the methods discussed. While modal synthesis stands out for its efficiency, it is solely applicable to linear models and necessitates pre-computation to determine the number of modes denoted as $N_{m}$. On the other hand, FDTD computes highly nonlinear and dynamic solutions but demands a substantial computational load due to iterative updates across both temporal ($N_{t}$) and spatial ($N_{x}$) samples. In contrast, differentiable audio processing methods offer efficient nonlinear sound synthesis, typically leveraging a smaller number of frames ($N_{r}$) compared to the total time samples ($N_{t}$). However, they often lack physical controllability. In this regard, we propose DMSP, which approximates the solution of Equation 3 efficiently by leveraging Equation 2 in the form of neural networks. Figure 1 provides a visual depiction of the DMSP.
(a) Network architecture of DMSP
(b) AM/FM Block
Figure 3: Network architecture. DMSP synthesizes a pitch skeleton with an inharmonic structure, drawing upon overtones derived from the modes of the string. The modes can either be derived directly using the modal decomposition (DMSP-Hybrid, the hybrid of DMSP and Modal), or using the neural network trained to estimate the modes (DMSP, the fully-neural-network method). Yet, relying solely on modal frequencies and corresponding shape functions delineates a linear solution, which falls short of capturing the nuances of nonlinear string motion. To address this, DMSP introduces FM and AM blocks to modulate the modes of the linear solution. This modulation process enables DMSP to estimate the pitch skeleton of the nonlinear solution. Consequently, the output waveform is synthesized through the spectral modeling pipeline, incorporating both (in)harmonic components and the filtered noise.
3.1 Problem Statement
The objective of this study is to establish a mapping from the parameter space to the solution space, utilizing a finite set of observations comprising parameter-solution pairs from this mapping. We delineate this objective as follows: Consider the partial differential equation depicted in Equation 3, applicable for $(x,t)$ within $\Omega\times[0,\infty)$, with clamped boundary conditions as specified in Equation 4, where $\Omega$ represents a bounded domain in $\mathbb{R}^{N_{x}}$. We assume that the solution $u:\Omega\times[0,\infty)\to\mathbb{R}$ resides within the Banach space $\mathcal{U}$. For a given PDE parameter $\rho\in\mathcal{P}$ and initial condition $u_{0}\in\mathcal{U}$, let $\mathcal{S}:\mathcal{P}\to\mathcal{U}$ denote a nonlinear map, specifically, the FDTD numerical solver tailored to the context of this study. Assume that we are provided with observations $\{\rho^{(i)},u^{(i)}\}_{i=1}^{N}$, where $\rho^{(i)}$ comprises independent and identically distributed (i.i.d.) samples drawn from a probability measure supported on $\mathcal{P}$, and $u^{(i)}=\mathcal{S}(\rho^{(i)})$ potentially contains noise. Our goal is to construct an approximation of $\mathcal{S}$ denoted as $\mathcal{S}_{\theta}:\mathcal{P}\to\mathcal{U}$, and select parameters $\theta^{*}\in\mathbb{R}^{N_{\theta}}$ such that $\mathcal{S}_{\theta^{*}}\approx\mathcal{S}$. Leveraging $\mathcal{S}_{\theta}$, one can compute the solution $\hat{u}=\mathcal{S}_{\theta}(\rho)$ corresponding to a new parameter $\rho\in\mathcal{P}$. By specifying values for $x$ and $t$, one can then either synthesize the sound of the string picked-up (also referred to as read-out) at a specific location $x_{0}$ as $\hat{u}(x_{0},t)$, or simulate the motion of the string by concatenating $\hat{u}(x,t)$ across all $x\in\Omega$. In practice, $\Omega$ and $[0,\infty)$ are bounded and discretized to form an evenly distributed spatio-temporal grid as $\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{t}}\subset\Omega\times[0,\infty)$, where the spatial grid samples are uniformly distributed in space according to the length of the equally spaced intervals $L/N_{x}$ and the temporal grid samples uniformly distributed according to the frequency of a fixed audio sampling rate.
3.2 Network Architecture
Parameter Encoder. To effectively capture material features inherent in the PDE parameter values, the parameter encoder leverages a random Fourier feature (RFF) layer [48, 49]. Given T60 frequencies $f_{\mathrm{T60}}^{(i)}$ and their corresponding times $t_{\mathrm{T60}}^{(i)}$ for $i=1,2$, the frequency-dependent damping coefficients $\sigma_{0}$ (and $\sigma_{1}$, if applicable) are derived using Equation 24. Subsequently, the frequency-independent damping factor $\exp(-\sigma_{0}t)$ is computed, explicitly multiplied by the mode amplitudes. All PDE parameters $\rho=\left\{\kappa,\alpha,\sigma_{0},\sigma_{1}\right\}\in\mathbb{R}^{4}\subset\mathcal{P}$ are encoded into a feature vector $h\in\mathbb{R}^{4\times d}$ with a Fourier embedding size of $d=256$.
AM and FM Blocks. As illustrated in Figure 3, DMSP employs amplitude modulation (AM) and frequency modulation (FM) modules to modulate the mode frequencies and amplitudes of the linear solution, as depicted in Equation 2, to synthesize the solution of the nonlinear wave described in Equation 3. We utilize two multilayer perceptrons (MLPs) for the modulation layers. Although a simpler and perhaps more conventional choice of architecture would be a GRU, to the decoder architecture of DDSP [16], we choose MLPs for their slightly better empirical results. It’s noteworthy that DDSP decodes the sinusoidal frequency envelope with fixed frequency values. In contrast, DMSP decodes both the envelope and the frequency values independently, employing two distinct MLP blocks, namely AM and FM.
Mode Estimator. As detailed in subsection 2.1, determining allowed values for mode frequencies and amplitudes, corresponding to specific initial and boundary conditions of the string, typically involves a root-finding process conducted offline. While these numerical solvers offer high accuracy up to a specified iterative threshold, they necessitate pre-computation, as illustrated in Table 1. The mode estimator module within DMSP estimates the modes from the initial condition using an MLP. The initial condition is parameterized by a pluck position $p_{x}$ and its peak amplitude $p_{a}$. Subsequently, the physical properties $\kappa$, $\gamma$, $\sigma_{0}$, and $\sigma_{1}$ are encoded using a random Fourier feature (RFF) layer. The mode frequencies and amplitudes are then estimated by the MLP, followed by the application of suitable scaling activations. It’s pertinent to note that the mode estimator operates independently and is trained separately from the other modules. During training, the ground truth modes (computed using the modal decomposition method) are fed into the AM and FM blocks, ensuring accurate synthesis while training the synthesis part of the model.
3.3 Loss Function
In training DMSP, we employ a combination of four loss terms: (1) waveform $\ell_{1}$ loss that measures the $L_{1}$ discrepancy between the synthesized waveform and the ground truth waveform, (2) Multi-scale spectral (MSS) loss that captures spectral differences across multiple scales, ensuring fidelity in spectral representation, and (3) Pitch loss ($\mathcal{L}_{f_{0}}$) that penalizes deviations from the ground truth fundamental frequency ($f_{0}$). Optimizing the frequency parameters of sinusoidal oscillators via gradient descent over an audio loss function poses a challenge due to the non-convex nature of the optimization problem, as highlighted in various studies [50, 51]. Damped sinusoids offer a workaround for the issue of non-convexity concerning frequency parameters [51], or alternative metrics are proposed to mitigate the risk of falling into bad local minima [52, 53]. In our approach, we adopt a parameter regression joint training strategy, akin to the pre-training phase of the work by Engel et al. [54]. Among the output mode frequencies, we optimize one frequency component to match the target fundamental frequency ($f_{0}$) annotated using CREPE [55]. As the comparison models are all aware of the frequency of the mode they are outputting, the models are updated to minimize the discrepancy between the fundamental frequency estimated as the lowest one of the modes (denoted by $\hat{f}_{0}$) and the ground-truth $f_{0}$ of the FDTD-simulated audio estimated via CREPE as $\mathcal{L}_{f_{0}}=\|\hat{f}_{0}-f_{0}\|_{1}$.
4 Experiments
This section introduces the empirical investigation that seeks to find the answers to the following questions: (1) how much does the modal synthesis-based inductive bias of a neural network advantage physical modeling sound synthesis, (2) how much does it improve nonlinearity and what physical analysis is possible, and (3) to what extent are techniques required for training such model.
4.1 Experimental Setup
Dataset. We use the StringFDTD-Torch [43], the open-source nonlinear string simulator, to compute the solution of the Equation 3 in a temporal sampling rate of 48 kHz and a spatial sampling rate. The solution is upsampled to a spatial resolution of 256 using a bivariate spline approximation over a rectangular spatio-temporal mesh upto the 5th order degree. We simulate 10263 different strings by randomly augmenting the material properties, e.g., $\kappa$, $\alpha$, $\sigma_{0}$, and $\sigma_{1}$, with various plucking profiles $u_{0}$. The simulation results in a total amount of 729.8 hours of wave files that corresponds to the 1-sec string sounds picked up at 256 different positions for each string. For the test data, 715 strings are newly synthesized with the parameters sampled in i.i.d. The test data consists of 336 linear ($\alpha=1$) and 379 nonlinear ($\alpha>1$) strings, each of which has a 256 spatial grid size, resulting in approximately 50 hours of wave files. Table 4 specifies the range of the sampled PDE parameters.
Baselines. As the first attempt to tackle the task of neural audio synthesis from the physical properties, we compare our proposed model to three other models. We consider three baselines, namely: Modal and DDSPish, and DDSPish without FM. Modal synthesis is the linear wave solution as in Equation 1, where the modal frequencies and the shape functions are pre-computed. DDSPish is a neural network based on the harmonic plus noise model, similar to DDSP. Yet, this -ish suffix emphasizes that this model is different from the DDSP model [16]. DDSPish does not have a reverb module but instead adds frequency modulation, and most notably it has a parameter encoder that allows generating sounds from physical parameters. Please see Appendix D for more details of the baselines.
Evaluation Metrics. We report results on three metrics, signal-distortion-ratio (SDR), scale-invariant signal-distortion-ratio (SI-SDR) [56], multi-scale spectral (MSS) distance, and the pitch difference in Hz. The MSS metric for the evaluation was computed using the short-time Fourier transformation (STFT) with 513 number of frequency bins and the hop size of 256 samples. The Pitch metric is computed as the $\ell^{1}$ norm of the difference between $\hat{f}_{0}$ and $f_{0}$.
4.2 Results
Table 2: Synthesis Results
Model Linear ($\alpha=1$) Nonlinear ($\alpha>1$)
SI-SDR SDR MSS Pitch SI-SDR SDR MSS Pitch
(dB, $\uparrow$) (dB, $\uparrow$) (dB, $\downarrow$) (Hz, $\downarrow$) (dB, $\uparrow$) (dB, $\uparrow$) (dB, $\downarrow$) (Hz, $\downarrow$)
Modal –3.191 0.681 18.449 0.420 –16.611 –1.900 17.254 2.316
DDSPish –39.478 –2.598 11.047 5.518 –25.951 –2.102 9.745 3.306
DDSPish-XFM –46.609 –2.257 10.911 11.304 –46.858 –2.272 10.299 14.013
DMSP-Hybrid –2.844 1.496 12.525 0.792 15.670 16.455 4.772 1.027
DMSP –22.298 –2.000 12.504 1.717 –10.315 0.221 5.656 1.437
Differentiable Sound Synthesis. The efficacy of DMSP is studied as shown in Table 2. First, the modal synthesis method calculates a linear solution for a damped stiff string, and the mode for the linear test set is calculated offline. The discrepancy between the outcomes of modal synthesis and FDTD in the linear case can be attributed to the absence of frequency-dependent damping in Equation 1. In other words, the frequency information serves as an oracle for the Modal, as evidenced by its lowest pitch score, but the decay of amplitude over time is not sufficiently modeled, which is where the remaining models demonstrate superior performance. Considering the DDSPish models, the MSS is approximately 1.6 dB ahead of the DMSPs, but the difference can reach up to 44 dB in the case of the SI-SDR. It is worth noting that $\alpha$ is uniformly sampled within the range $(1,25)$ for the training data, falling into the category of the nonlinear strings. In the case of nonlinear test set, the difference between the modal solution and the FDTD solution becomes larger. On the other hand, DDSPish models based on spectral modeling synthesis but learned without using any modal information show the lowest performance. The superiority of DMSP is even more pronounced in the nonlinear case. While Modal, which can only cover the solution to the linear case, shows attenuated performance, the DMSPs demonstrate the best performance in all metrics. In particular, DMSP-Hybrid, which precomputes the mode frequency like Modal, performs FM and AM for the nonlinear solution, showing an SI-SDR improvement of nearly 31 dB and an MSS improvement of nearly 13 dB over Modal. DMSP, which estimates the mode as a neural network without the pre-computation step, also outperforms Modal on all metrics for the nonlinear case. The primary reason for the performance discrepancy between the nonlinear case and the linear case for DMSP is that the training data is rarely precisely equal to $1$ in the $\alpha$ distribution when it is sampled.
(a) $p_{x}=0.1$
(b) $p_{x}=0.2$
(c) $p_{x}=0.3$
(d) $p_{x}=0.4$
Figure 4: Visualization of the string displacement over time (horizontal) and space (vertical). For different initial conditions, the results synthesized by DMSP are shown as solid black lines and those simulated by FDTD as dashed gray lines.
Figure 5: Objective scores over physical parameters.
Controllable Physical Simulation. The quantitative scores for various physical condition parameters are visualized in Figure 5. Trends show how the results of Modal synthesis and DMSP vary for different pickup positions ($x$), stiffness ($\kappa$), tension ($\alpha$), pluck amplitude ($p_{a}$), and pluck position ($p_{x}$). Of these, $\alpha$ and $p_{a}$ in particular are known to increase the nonlinearity of the string as they increase in magnitude, which can be seen by the lower Modal synthesis scores. For DMSP, we see an overall improvement in the score, with a lower propensity for nonlinearity. Figure 4 depicts the simulated state of the string as the pluck position in the initial condition is varied. The results synthesized by DMSP can reconstruct a very accurate initial condition, similar to the results simulated by FDTD. The vibration propagating through time along the string exhibits a distinct behavior contingent upon the initial condition. FDTD employs a recursive calculation of displacement, necessitating a number of iterations equal to the number of samples at the audio sampling rate. In contrast, DMSP is capable of obtaining the desired displacement in both time and space simultaneously.
String Sound Synthesis. Spectrograms of the test samples are visualized in Figure 6. The instantaneous frequencies are identified in a rainbow color map, where the color intensities represent the logarithmic magnitude of the power spectra. The FDTD-simulated spectrogram clearly shows pitch glide and phantom partials at the beginning of the pluck. In contrast, modal synthesis methods that model linear solutions do not show these nonlinear characteristics. The DDSPish-XFM model employs a harmonic pitch skeleton comprising integer multiples of $f_{0}$, thereby precluding the inharmoicities resulting from stiffness. The DDSPish model demonstrates enhanced mode estimation capabilities through learned FM, which modulates the harmonic pitch skeleton to be inharmonic. However, there is scope for further improvement in frequency estimation, particularly in instances where the FM learning process is unstable for high frequencies. On the other hand, the DMSP model, which estimates the mode frequency and amplitude from $u_{0}$, shows an improved pitch skeleton and stable frequency estimation. The DMSP-Hybrid model, which learns to AM and FM the sinusoidal oscillators of the modal solution that requires mode precomputation, shows the most similar results to FDTD.
(a) Modal
(b) DDSPish-XFM
(c) DDSPish
(d) DMSP
(e) DMSP-Hybrid
(f) FDTD (GT)
Figure 6: Spectrogram of the synthesized samples on the test set.
Table 3: Ablation Study
Model Linear Nonlinear
SI-SDR SDR MSS Pitch SI-SDR SDR MSS Pitch
(dB, $\uparrow$) (dB, $\uparrow$) (dB, $\downarrow$) (Hz, $\downarrow$) (dB, $\uparrow$) (dB, $\uparrow$) (dB, $\downarrow$) (Hz, $\downarrow$)
DMSP-Hybrid –2.844 1.496 12.525 0.792 15.670 16.455 4.772 1.027
   w.o. $\mathcal{L}_{f_{0}}$ –2.919 0.774 13.487 0.792 –5.418 1.509 8.983 2.653
DMSP –22.298 –2.000 12.504 1.717 –10.315 0.221 5.656 1.437
   w.o. $\mathcal{L}_{f_{0}}$ –21.351 –2.699 13.482 1.717 –16.435 –1.074 9.060 2.922
Ablation Study. The ablation study on the choice of the training loss functions is presented in Table 3. Overall, for linear strings, the performance does not vary much depending on which loss is used, especially for pitch. This is due to the alpha masking applied to the FM block, which is designed to prevent FM from occurring when alpha is 1. This is the main factor that determines the nonlinearity of the string. For AM, there is no masking depending on alpha, so the remaining metrics except pitch do vary. For nonlinear data, the learning results vary depending on the loss design.
Figure 7: Simulated string state visualization.
Motion Synthesis. The main advantage of this neural net based on modal synthesis is that it can synthesize not only sound but also motion, which is one of the main characteristics of physical modeling techniques. In particular, the DMSP is able to visualize the corresponding string motion as a video, despite the fact that the receptive field and computational complexity required to obtain a solution $u(x,t)$ for a single spatio-temporal point is of order 1, as shown in Table 1, when these solutions are pooled for a given $x\in\Omega$ and $t\in[0,1)$, the corresponding string motion can be visualized as a video. Figure 7 visualizes the resulting transverse displacement of the string over time (horizontal axis) and space (vertical axis). As mentioned earlier, the transverse displacement of the FDTD is coupled to the longitudinal motion, which is why it differs from the Modal synthesis output, which synthesizes the motion of a linear damped stiff string. The results output by DMSP show improved accuracy.
5 Conclusion
We present a novel neural network-based method that efficiently simulates plucked string motions. Our differentiable modal synthesis for physical modeling (DMSP) can simulate a dynamic nonlinear string motion by synthesizing the sound using the physical properties of the string. It is an efficient approximation of existing physical modeling methods. We demonstrate the efficacy of training the neural network using mode frequency information by extending the DDSP with a modal synthesis pipeline. This opens the door to a new field of differentiable audio signal processing, extending it to the field of physical modeling for musical sound synthesis. While the proposed method offers control over several physical parameters of a musical instrument, it still faces limitations in terms of generalizing to physical parameters and sounds in real-world measurements. This study paves the way for future research in this area. To the best of our knowledge, this is the first study to simultaneously synthesize sound and motion from the properties of a stringed instrument.
Appendix A Proof of the Linear Damped Stiff String Solution
Proposition 1.
A solution to the damped linear stiff string model Equation 1 with a clamped boundary condition $u(\pm L/2,t)=u_{x}(\pm L/2,t)=0$ and the initial condition given as $u(x,0)=u_{0}(x)$ and $\partial_{t}u(x,0)=0$ can be expressed as
$\begin{split}u(x,t)&=\sum_{n=1}^{\infty}X_{n}(x)T_{n}(t)\\ X_{n}(x)&=c_{1}\left(\sin\mu_{n}x-\frac{\sin\mu_{n}L/2}{\sinh\nu_{n}L/2}\sinh\nu_{n}x\right)+c_{2}\left(\cos\mu_{n}x-\frac{\cos\mu_{n}L/2}{\cosh\nu_{n}L/2}\cosh\nu_{n}x\right)\\ T_{n}(t)&=e^{-\sigma_{0}t}\cos\left(\sqrt{\mu_{n}^{4}\kappa^{2}+\mu_{n}^{2}\gamma^{2}-\sigma_{0}^{2}}\right)t\end{split}$
Proof.
As it is hinted in Equation 2a, the procedure to derive Equation 2b and Equation 2c uses the method of separation of variables. The derivation consists of three main steps that start by trying the ansatz $u(x,t)=X(x)T(t)$. Substituting this ansatz into the Equation 1 gives
$\gamma^{2}\frac{X^{\prime\prime}}{X}-\kappa^{2}\frac{X^{(4)}}{X}=\frac{T^{\prime\prime}}{T}+2\sigma_{0}\frac{T^{\prime}}{T}=\begin{cases}\varsigma\\ 0\\ -\varsigma\\ \end{cases}$ (5)
with $\varsigma\in\mathbb{R}$ as the separation constant.
  1. 1.
    Solving for $T$.
    $T^{\prime\prime}+2\sigma_{0}T\pm\varsigma^{2}T=0$ (6)
    Roots of the characteristic polynomial of this equation are $\beta_{\pm}=-\sigma_{0}\pm\sqrt{\sigma_{0}^{2}\mp\varsigma^{2}}$. Three solutions are available:
    • Overdamping ($\sigma_{0}^{2}>\varsigma^{2}$): $T=A_{1}e^{\beta_{+}t}+A_{2}e^{\beta_{-}t}$
    • Critical damping ($\sigma_{0}^{2}=\varsigma^{2}$): $T=(A_{1}+A_{2}t)e^{-\sigma_{0}t}$
    • Underdamping ($\sigma_{0}^{2}<\varsigma^{2}$): Rewrite the roots as $\beta_{+}=-\sigma_{0}+i\omega$ and $\beta_{-}=-\sigma_{0}-\hat{\omega}$ with $\omega=\sqrt{\varsigma^{2}-\sigma_{0}^{2}},\hat{\omega}=\sqrt{\varsigma^{2}+\sigma_{0}^{2}}$, then
      $T=e^{-\sigma_{0}t}(A_{1}e^{i\omega t}+A_{2}e^{-\hat{\omega}t})$ (7)
      The initial condition $\partial_{t}u(x,0)=0$ implies $A_{2}=0$ yielding the real solution of the form $T=A_{1}e^{-\sigma_{0}t}\cos\omega t$. This implies that the only valid root is $\beta_{+}=-\sigma_{0}+i\omega$, hence it is enough to consider the $-\varsigma$ case only, in the Equation 5.
    This work only considers the underdamped case where the $\sigma_{0}$ is sufficiently small, as it should be in a reasonable string model for any musical purposes.
  2. 2.
    Solving for $X$.
    $X^{(4)}-\frac{\gamma^{2}}{\kappa^{2}}X^{\prime\prime}-\frac{\varsigma^{2}}{\kappa^{2}}X=0$ (8)
    Substitute $l=\gamma^{2}/2\kappa^{2}$ and $m=\varsigma^{2}/\kappa^{2}$. The roots of the characteristic polynomial for this equation are ±$\pm\sqrt{l\pm\sqrt{l^{2}+m}}$ where $l-\sqrt{l^{2}+m}\leq 0$. Therefore the solution is rewritten as
    $X(x)=\underbrace{B_{1}\sin\mu x+B_{2}\sinh\nu x}_{\text{odd function}}+\underbrace{B_{3}\cos\mu x+B_{4}\cosh\nu x}_{\text{even function}}$ (9)
    where $\mu=\sqrt{\sqrt{l^{2}+m}-l}$ and $\nu=\sqrt{\sqrt{l^{2}+m}+l}$. By applying the boundary condition $u(L/2,t)=0$ to the even and odd functions each, then obtain
    $B_{2}=-\frac{\sin\mu\frac{L}{2}}{\sinh\nu\frac{L}{2}}B_{1}\quad\text{and}\quad B_{4}=-\frac{\cos\mu\frac{L}{2}}{\cosh\nu\frac{L}{2}}B_{3}$ (10)
    reducing the solution as
    $X(x)=B_{1}\left(\sin\mu x-\frac{\sin\mu\frac{L}{2}}{\sinh\nu\frac{L}{2}}\sinh\nu x\right)+B_{3}\left(\cos\mu x-\frac{\cos\mu\frac{L}{2}}{\cosh\nu\frac{L}{2}}\cosh\nu x\right).$ (11)
  3. 3.
    Finding allowed values that satisfy the conditions.
    Substituting the ansatz by Equation 7 and Equation 11 summarizes the solution as
    $\begin{split}u(x,t)&=e^{-\sigma_{0}t}\cos\left(\sqrt{\mu^{4}\kappa^{2}+\mu^{2}\gamma^{2}-\sigma_{0}^{2}}\cdot t\right)\\ &\times\Bigg{[}c_{1}\left(\sin\mu x-\frac{\sin\mu\frac{L}{2}}{\sinh\nu\frac{L}{2}}\sinh\nu x\right)+c_{2}\left(\cos\mu x-\frac{\cos\mu\frac{L}{2}}{\cosh\nu\frac{L}{2}}\cosh\nu x\right)\Bigg{]}\end{split}$ (12)
    As mentioned earlier, the coefficients $c_{1}=A_{1}B_{1}$ and $c_{2}=A_{1}B_{3}$ are determined by the allowed values that satisfy the initial condition $\sum_{n=1}^{\infty}X_{n}(x)=u_{0}(x)$. To determine the allowed values of $\mu$, apply the boundary condition $u(L/2,t)=u_{x}(L/2,t)=0$ to the even function that gives
    $B_{3}\cos\mu\frac{L}{2}=-B_{4}\cosh\nu\frac{L}{2}\quad\text{and}\quad-\mu B_{3}\sin\mu\frac{L}{2}=-\nu B_{4}\sinh\nu\frac{L}{2}$ (13)
    or alternatively,
    $\mu\tan\mu\frac{L}{2}=-\nu\tanh\nu\frac{L}{2}.$ (14)
    Similarly, applying the same boundary condition to the odd function gives
    $B_{3}\sin\mu\frac{L}{2}=-B_{4}\sinh\nu\frac{L}{2}\quad\text{and}\quad\mu B_{3}\cos\mu\frac{L}{2}=-\nu B_{4}\cosh\nu\frac{L}{2}$ (15)
    or equivalently,
    $\nu\tan\mu\frac{L}{2}=\mu\tanh\nu\frac{L}{2}.$ (16)
    By finding the values of $\mu$ and $\nu=\sqrt{\mu^{2}+2l}$ that satisfy the Equation 14 and Equation 16 as $\mu_{n}$ and $\nu_{n}$, one can determine the allowed values of $m_{n}=(\mu_{n}^{2}+l)^{2}-l^{2}$ and $\varsigma_{n}=\sqrt{m_{n}\kappa^{2}}$.
Following these three steps gives the $n$-th mode oscillation of $u(x,t)$ as $X_{n}(x)T_{n}(t)$ where
$X_{n}(x)=c_{1}\left(\sin\mu_{n}x-\frac{\sin\mu_{n}L/2}{\sinh\nu_{n}L/2}\sinh\nu_{n}x\right)+c_{2}\left(\cos\mu_{n}x-\frac{\cos\mu_{n}L/2}{\cosh\nu_{n}L/2}\cosh\nu_{n}x\right)$ (17)
and
$T_{n}(t)=e^{-\sigma_{0}t}\cos\left(\sqrt{\mu_{n}^{4}\kappa^{2}+\mu_{n}^{2}\gamma^{2}-\sigma_{0}^{2}}\right)t.$ (18)
Hence, the analytic solution is expressed as a superposition of modes, where the infinite summation as in Equation 2a finalizes the modal expression of the analytic solution. ∎
Appendix B Nonlinear Damped Stiff String Vibration
The Kirchhoff–Carrier model stands out as the simplest representation of nonlinear distributed strings. It effectively captures the significant pitch glide effect and can be easily simulated using relatively straightforward finite difference schemes. However, it’s important to note that in this model, only the transverse motion is explicitly accounted for, while the longitudinal motion is averaged across the string’s length. Although this approximation may suffice under specific conditions, the interaction between longitudinal and transverse motion results in perceptually crucial effects like phantom partials. Hence, a more intricate model is required to encompass these phenomena. A comprehensive model of string vibration, which incorporates both longitudinal and transverse motion within a single plane (in dimensional form), is outlined as follows:
$\rho A\partial_{tt}u$ $=EA\partial_{xx}u-(EA-T_{0})\partial_{x}\left(\frac{\partial\Phi}{\partial(\partial_{x}u)}\right)$ (19a)
$\rho A\partial_{tt}\zeta$ $=EA\partial_{xx}\zeta-(EA-T_{0})\partial_{x}\left(\frac{\partial\Phi}{\partial(\partial_{x}\zeta)}\right)$ (19b)
The constants $\rho$ and $T_{0}$ are the material density and tension, respectively. In order to model the stiffness, Young’s modulus $E$ and the cross-sectional area $A$ are introduced. The function $\Phi$, which nonlinearly connects the two equations, is defined as follows:
$\Phi=\sqrt{(1+\zeta_{x})^{2}+u_{x}^{2}}-1-\zeta_{x}$ (20)
Note that the final term $-1-\zeta_{x}$ does not affect the dynamics of system Equation 19; it is included solely to adjust the zero-point energy of the entire system. Substituting $x^{\prime}=x/L$, $u^{\prime}=u/L$, and $\zeta^{\prime}=\zeta/L$ into the above system, and then removing the primes, one obtains:
$\partial_{tt}u$ $=\gamma^{2}\partial_{xx}u-\gamma^{2}(\alpha^{2}-1)\partial_{x}\left(\frac{\partial\Phi}{\partial q}\right)$ (21a)
$\partial_{tt}\zeta$ $=\gamma^{2}\alpha^{2}\partial_{xx}\zeta-\gamma^{2}(\alpha^{2}-1)\partial_{x}\left(\frac{\partial\Phi}{\partial p}\right)$ (21b)
where the parameters $\gamma$ and $\alpha$ are defined by
$\gamma=\frac{1}{L}\sqrt{\frac{T_{0}}{\rho A}}\qquad\text{and}\qquad\alpha=\sqrt{\frac{EA}{T_{0}}}.$ (22)
For $\Phi$, one has $\Phi(p,q)=\sqrt{(1+p)^{2}+q^{2}}-1-p$. Series approximations have been instrumental in analyzing nonlinear systems like the string model mentioned earlier; approximations up to third or fourth order are frequently used [35]. The function $\Phi(p,q)$ can be approximated with a variety of orders in $p$ and $q$:
$\Phi_{2}=\frac{1}{2}q^{2}\qquad\Phi_{3}=\frac{1}{2}q^{2}-\frac{1}{2}pq^{2}\qquad\Phi_{4}=\frac{1}{2}q^{2}-\frac{1}{2}pq^{2}+\frac{1}{2}q^{2}p^{2}-\frac{1}{8}q^{4}$
In this paper, we consider the following approximation to $\Phi(p,q)$, similar to $\Phi_{4}$, but lacking one of the fourth-order terms, following Bilbao [10].
$\Phi_{4}^{*}=\frac{1}{2}q^{2}-\frac{1}{2}pq^{2}-\frac{1}{8}q^{4}$
Under this choice $\Phi_{4}^{*}$, the system Equation 21 reduces to
$\partial_{tt}u$ $=\gamma^{2}\alpha^{2}\partial_{xx}u-\gamma^{2}(\alpha^{2}-1)\partial_{x}\left(q^{3}+2pq\right)$ (23a)
$\partial_{tt}\zeta$ $=\gamma^{2}\alpha^{2}\partial_{xx}\zeta-\gamma^{2}(\alpha^{2}-1)\partial_{x}\left(q^{2}\right)$ (23b)
which can be augmented by the stiffness and the damping term to become Equation 3. The nonlinear damped stiff string system as Equation 3 is very similar to the work by Bank and Sujbert [57] but with the damping coefficients given as the same values for the longitudinal and transverse directions.
Appendix C Damping coefficient
The damping coefficients $\sigma_{0}$ and $\sigma_{1}$ are typically determined experimentally, especially those for nonlinear systems. However, it can be somewhat difficult to set the values without any prior knowledge of the observations. In this regard, the practical settings for the damping coefficients can be approximately derived from the decay times, by resorting to the values for linear systems. The linear string damping coefficients $\sigma_{0}$ and $\sigma_{1}$ are derived from the (frequency-dependent) T60 values: where $f_{\mathrm{T60}}^{(1)}$ and $f_{\mathrm{T60}}^{(2)}$ denotes the two distinct frequencies, and $t_{\mathrm{T60}}^{(1)}$ and $t_{\mathrm{T60}}^{(2)}$ denotes the corresponding decay times. In this case,
$\sigma_{0}$ $=\frac{6\log(10)}{\xi_{1}-\xi_{2}}\left(\frac{\xi_{1}}{t_{\mathrm{T60}}^{(2)}}-\frac{\xi_{2}}{t_{\mathrm{T60}}^{(1)}}\right)$ (24a)
$\sigma_{1}$ $=\frac{6\log(10)}{\xi_{1}-\xi_{2}}\left(\frac{1}{t_{\mathrm{T60}}^{(1)}}-\frac{1}{t_{\mathrm{T60}}^{(2)}}\right)$ (24b)
where $\xi_{i}=-\gamma^{2}+\sqrt{\gamma^{4}+4\kappa^{2}\times(2\pi f_{\mathrm{T60}}^{(i)})^{2}}$ for $i=1,2$. For this model, as the energy loss increases monotonically with frequency, one must choose $t_{\mathrm{T60}}^{(1)}\geq t_{\mathrm{T60}}^{(2)}$ when $f_{\mathrm{T60}}^{(1)}<f_{\mathrm{T60}}^{(2)}$.
Figure 8: The DDSPish is designed in a similar way to DMSP,
Appendix D Baselines
In this section, we present more details of the baselines. All models, including DMSP, are trained using RAdam optimizer, with Noam learning rate scheduler with a peak learning rate of $10^{-3}$ reaching at 1000 number of warmup steps. Figure 8 summarizes the architectures of DDSPish models. The only architectural difference between DDSPish and DDSPish-XFM is the existence of the FM module. DDSPish-XFM is harmonic as the mode frequencies are not modulated after its initialization by the integer multiples of $f_{0}$, while DDSPish is inharmonic as FM is trained to modulate the harmonic pitch skeleton to match the inharmonic mode frequencies.
For the Modal synthesis, we compute the solution Equation 1 with the allowed values of $\mu_{n}$, $\nu_{n}$, $c_{1}$, and $c_{2}$ obtained using the Levenberg–Marquardt algorithm. We compute the modes up to the 100th order with the double-precision floating-point arithmetics. The modes are then post-processed to be cut under the Nyquist limit. Subsequently, all models are trained using 40 numbers of modes. We use 65 number of bands used for the filtered noise.
Table 4: PDE parameter sampling range
Min. Max. Unit
$f_{0}$ 98.00 440.0 Hz
$\kappa$ $0.01\gamma$ $0.03\gamma$
$\alpha$ 1 25 -
$t_{\mathrm{T60}}^{(1)}$ 10 25 sec
$t_{\mathrm{T60}}^{(2)}$ 10 30 sec
$f_{\mathrm{T60}}^{(1)}$ 1100 1200 Hz
$f_{\mathrm{T60}}^{(2)}$ 100 $f_{\mathrm{T60}}^{(1)}-1000$ Hz
$p_{a}$ 0.001 0.02 -
$p_{x}$ 0.1 0.5 -
Appendix E Datasets
As described in subsection 4.1, this paper utilizes the nonlinear string simulator, StringFDTD-Torch, presented by Lee et al. [43]. For the simulation, PDE parameters are uniformly random-sampled within the moderate parameter ranges. Table 4 summarizes the infimum and the supremum values for each uniform distribution of the PDE parameter. For a given random-sampled parameter set, the simulator outputs the transverse and the longitudinal solutions ($u$ and $\zeta$, resp.) while this work considers $u$ only. Yet, it is worth mentioning that the adopted $u$ is different from that of the one-dimensional string since $u$ and $\zeta$ are coupled as evident in Equation 3. The obtained $u$ is defined over a spatio-temporal grid, where the spatial resolutions are carefully chosen to mitigate the numerical stability criteria and the numerical dispersions while keeping the temporal resolution by the prefixed audio sampling rate. As the grid spacing is fixed by these values, the raw simulation data are consisted of diverse grid sizes depending on $f_{0}$, $\kappa$, and $\alpha$, making it difficult to batchify for training. For this reason, we spatially upsample the data to a fixed spatial grid size of 256, using a bivariate spline approximation over a rectangular spatio-temporal mesh up to the 5th order degree.